# Third quantization with Hartree approximation for open-system bosonic transport

### v2

@@ I Introduction
The emergence of ultracold-atoms as a versatile platform for studying quantum physics in and out of equilibrium has brought many interesting simulations and analogues of condensed matter physics [1, 2, 3, 4]. For example, quantum transport has been an important subject in electronic systems [5, 6], and it has received attention in atomic systems as well [7]. While bosonic and fermionic atoms have been used in experiments, the bosonic physics has less counterparts in conventional condensed matter physics and poses challenges for theoretical understanding because of the infinite-dimensional Fock space allowed by the Bose-Einstein statistics. Therefore, some methods developed for electronic transport [8, 9] may not apply to bosonic systems directly.
@@
Nevertheless, interactions in the system result in many interesting transport phenomena, including diode effect [14, 15, 39] and conducting-non conducting transition [40]. Tensor-based methods have been a useful tool for studying 1D interacting systems, including up to 200 sites with 4 bosonic states per site of the 1D BHM [41] and up to 96 sites and 6 bosons per site of a two-species Bose-Hubbard ring [42]. A combination of the mean-field theory and density-matrix renormalization group method allows the study of the correlation functions for large (up to 3700-site) systems with a cap of three bosons per site [43]. However, the Fock space of bosons has been truncated to render the calculations manageable. Here we introduce Hartree approximation [44, 45, 46, 47, 48] to the third quantization of bosons without truncating the bosonic Fock space by approximating the interaction term of the BHM with quadratic terms. Utilizing the analytic framework of third quantization provides an efficient and scalable method to find the steady-state solution of the driven BHM, described by the Lindblad equation, without truncating the infinite Fock space of bosons.
We will examine three examples of the BHM: the uniform chain, interaction induced diode effect, and Su-Schrieffer-Heeger (SSH) Hubbard model. The first serves as a calibration of the results, the second has been studied in the context of atomtronics [14, 15], and the third is based on the SSH model with alternating hopping coefficients [49, 50], which has illustrated interaction effects in topological systems [51, 52, 53, 54]. By comparing the results of small systems with the exact interaction term but in finite truncated Fock space, the 3rd Q with Hartree approximation correctly captures the qualitative features in all cases. However, it over-estimates the steady-state current due to the mean-field treatment of the interactions. Nevertheless, we are able to study the steady states of bosonic systems having thousands of lattice sites without truncating the Fock space, allowing an interpolation between small and large systems. The 3rd Q with Hartree approximation thus serves as a fast and scalable estimation of the upper-bound of bosonic transport for large systems with many bosons in the weak interaction regime and complements other available numerical methods.
@@ III.3 Bosonic SSH-Hubbard model
As discussed in Ref. [69] for the fermionic SSH model, localized edge states from topological systems can couple to the reservoirs and affect quantum transport. In the present case of bosonic transport, the two ends of the SSH-Hubbard model have either both $t_{1}$ or both $t_{2}$ connected to the reservoirs. The different couplings between the edges of the system and the reservoirs in the forward and reverse configurations due to the swap of $t_{1,2}$ thus lead to the rectification when we compare the currents through different configurations. This argument is also supported by the aforementioned observation that $U$ in general enhances $R$ as shown in Fig. 7 (a) and the argument that interaction induces scattering effects regardless of the hopping coefficients, thereby reducing any difference from swapping the configurations. To further confirm the argument about system-reservoir coupling, we show in Fig. 7 (b) the rectification ratio as a function of $t_{2}/t_{1}$ for the noninteracting ($U=0$) SSH model with different values of the system-reservoir coupling $\gamma$. Indeed, $R$ increases with $\gamma$ because the system-reservoir coupling determines how efficient particles can be exchanged. A higher value of $\gamma$ implies a higher exchange between the system and reservoirs regardless of the hopping patterns, so the results are less sensitive to the ratio $t_{2}/t_{1}$. Moreover, the noninteracting SSH model approaches perfect rectification ($R=0$) as $t_{2}/t_{1}\rightarrow 0$ when $\gamma\rightarrow 0$, but in that limit the current also approaches zero in both configurations due to the vanishing system-reservoir coupling. A similar increase of $R$ with $\gamma$ for the case with $U/t_{1}=1$ is shown in Fig. 7 (c), so the argument works for both noninteracting and interacting systems.
@@ IV Discussion
The three examples studied here are realizable in cold-atom or condensed-matter experiments. For example, Ref. [75] demonstrates the 1D BHM with uniform or tilted parameters by loading atoms into optical lattices. It was also used to measure energy ensemble averages in optical lattices [76]. The atomtronic battery has been realized via a double-well  [18], and diode-like behavior have been studied in a similar setup with an adjoined well [19]. In addition to atomtronic diode effect [14, 15], diode effect may also be engineered with transmission line resonators and superconducting qubits [77]. Nevertheless, our examples of diode effect in the BHM with inhomogeneous interactions or alternating hopping coefficients show that many tunable parameters can be used to induce asymmetric currents in quantum bosonic systems as the configuration changes. The appearance of diode effect in electronic and bosonic systems shows that it transcends spin statistics and will continue to be an interesting research topic in condensed-matter and atomic physics.

### v3

@@ I Introduction
The emergence of ultracold-atoms as a versatile platform for studying quantum physics in and out of equilibrium has brought many interesting simulations and analogues of condensed matter physics [1, 2, 3, 4]. For example, quantum transport has been an important subject in electronic systems [5, 6], and it has received attention in atomic systems as well [7]. While bosonic and fermionic atoms have been used in experiments, bosonic physics has less counterparts in conventional condensed matter physics and poses challenges for theoretical understanding because of the infinite-dimensional Fock space allowed by the Bose-Einstein statistics. Therefore, some methods developed for electronic transport [8, 9] may not apply to bosonic systems directly.
@@
Nevertheless, interactions in the system result in many interesting transport phenomena, including diode effect [14, 15, 39] and conducting-non conducting transition [40]. Tensor-based methods have been a useful tool for studying 1D interacting systems, including up to 200 sites with 4 bosonic states per site of the 1D BHM [41] and up to 96 sites and 6 bosons per site of a two-species Bose-Hubbard ring [42]. A combination of the mean-field theory and density-matrix renormalization group method allows the study of the correlation functions for large (up to 3700-site) systems with a cap of three bosons per site [43]. However, the Fock space of bosons has been truncated to render the calculations manageable. Here we introduce the Hartree approximation [44, 45, 46, 47, 48] to the third quantization of bosons without truncating the bosonic Fock space by approximating the interaction term of the BHM with quadratic terms. Utilizing the analytic framework of third quantization provides an efficient and scalable method to find the steady-state solution of the driven BHM, described by the Lindblad equation, without truncating the infinite Fock space of bosons. We mention that multi-configurational Hartree (or Hartree-Fock) approximations have been implemented in many-body calculations [49, 50, 51, 52, 53, 54], and there are other approaches, such as phase-space methods using the Wigner function [55, 56, 57, 58, 59, 60, 61], to study interacting systems in or out of equilibrium.
We will examine three examples of the BHM: the uniform chain, interaction-induced diode effect, and Su-Schrieffer-Heeger (SSH) Hubbard model. The first serves as a calibration of the results, the second has been studied in the context of atomtronics [14, 15], and the third is based on the SSH model with alternating hopping coefficients [62, 63], which has illustrated interaction effects in topological systems [64, 65, 66, 67]. By comparing the results of small systems with the exact interaction term but in finite truncated Fock space, the 3rd Q with Hartree approximation correctly captures the qualitative features in all cases. However, it over-estimates the steady-state current due to the mean-field treatment of the interactions. Nevertheless, we are able to study the steady states of bosonic systems having thousands of lattice sites without truncating the Fock space, allowing an interpolation between small and large systems. The 3rd Q with Hartree approximation thus serves as a fast and scalable estimation of the upper-bound of bosonic transport for large systems with many bosons in the weak interaction regime and complements other available numerical methods.
@@ III.3 Bosonic SSH-Hubbard model
As discussed in Ref. [82] for the fermionic SSH model, localized edge states from topological systems can couple to the reservoirs and affect quantum transport. In the present case of bosonic transport, the two ends of the SSH-Hubbard model have both $t_{1}$ or both $t_{2}$ connected to the reservoirs. The different couplings between the edges of the system and the reservoirs in the forward and reverse configurations due to the swap of $t_{1,2}$ thus lead to the rectification when we compare the currents through different configurations. This argument is also supported by the aforementioned observation that $U$ in general enhances $R$ as shown in Fig. 7 (a) and the argument that interaction induces scattering effects regardless of the hopping coefficients, thereby reducing any difference from swapping the configurations.
To further confirm the argument about system-reservoir coupling, we show in Fig. 7 (b) the rectification ratio as a function of $t_{2}/t_{1}$ for the noninteracting ($U=0$) SSH model with different values of the system-reservoir coupling $\gamma$. Indeed, $R$ increases with $\gamma$ because the system-reservoir coupling determines how efficient particles can be exchanged. A higher value of $\gamma$ implies more frequent exchanges between the system and reservoirs regardless of the hopping patterns, so the results are less sensitive to the ratio $t_{2}/t_{1}$. Moreover, the noninteracting SSH model approaches perfect rectification ($R=0$) as $t_{2}/t_{1}\rightarrow 0$ when $\gamma\rightarrow 0$, but in that limit the current also approaches zero in both configurations due to the vanishing system-reservoir coupling. A similar increase of $R$ with $\gamma$ for the case with $U/t_{1}=1$ is shown in Fig. 7 (c), so the argument works for both noninteracting and interacting systems.
@@ IV Discussion
The three examples studied here are realizable in cold-atom or condensed-matter experiments. For example, Ref. [88] demonstrates the 1D BHM with uniform or tilted parameters by loading atoms into optical lattices. It was also used to measure energy ensemble averages in optical lattices [89]. The atomtronic battery has been realized via a double-well  [18], and diode-like behavior has been studied in a similar setup with an adjoined well [19]. In addition to atomtronic diode effect [14, 15], diode effect may also be engineered with transmission line resonators and superconducting qubits [90]. Nevertheless, our examples of the diode effect in the BHM with inhomogeneous interactions or alternating hopping coefficients show that many tunable parameters can be used to induce asymmetric currents in quantum bosonic systems as the configuration changes. The appearance of diode effect in electronic and bosonic systems shows that it transcends spin statistics and will continue to be an interesting research topic in condensed-matter and atomic physics.
@@
The 3rd Q with Hartree approximation also complements other many-body methods. For example, multi-configurational Hartree (or Hartree-Fock) approximations [49, 50, 51, 52, 53, 54] mainly focus on finite-time dynamics of ground-state properties while the 3rd Q with Hartree approximation can reveal the steady-state behavior in the long-time limit of general pure or mixed states. Meanwhile, phase-space methods [55, 56, 57, 58, 59, 60, 61] are built on kinetic equations or path-integral formalisms whereas the 3rd Q with Hartree approximation is based on the Lindblad equation common in the open quantum system literature.